### v1

# XRF V2: A Dataset for Action Summarization with Wi-Fi Signals, and IMUs in Phones, Watches, Earbuds, and Glasses

Abstract.
Human Action Recognition (HAR) plays a crucial role in applications such as health monitoring, smart home automation, and human-computer interaction. While HAR has been extensively studied, action summarization, which involves identifying and summarizing continuous actions, remains an emerging task. This paper introduces the novel XRF V2 dataset, designed for indoor daily activity Temporal Action Localization (TAL) and action summarization. XRF V2 integrates multimodal data from Wi-Fi signals, IMU sensors (smartphones, smartwatches, headphones, and smart glasses), and synchronized video recordings, offering a diverse collection of indoor activities from 16 volunteers across three distinct environments. To tackle TAL and action summarization, we propose the XRFMamba neural network, which excels at capturing long-term dependencies in untrimmed sensory sequences and outperforms state-of-the-art methods, such as ActionFormer and WiFiTAD. We envision XRF V2 as a valuable resource for advancing research in human action localization, action forecasting, pose estimation, multimodal foundation models pre-training, synthetic data generation, and more.
dataset, continuous action recognition, temporal action localization, action summarization, Mamba, wearable devices, smartphone, smartwatch, glasses, earbuds, IMU, Wi-Fi, LLM Agent, smart home, ambient sensing
copyright: acmlicensedjournalyear: 2025publicationmonth: 1doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Ubiquitous and mobile computing systems and tools
1. Introduction
Human Indoor Action Understanding (HIAU) is pivotal in enabling intelligent environments, with wide-ranging applications in health monitoring, smart home automation, and human-computer interaction. Currently, human action recognition (HAR) is the most extensively studied aspect in HIAU. HAR aims to identify and categorize discrete human actions by leveraging data from diverse modal sensors, such as smartwatches (Laput and Harrison, 2019; Wang et al., 2024b), smartphones (Sun et al., 2024; DeVrio et al., 2023), earbuds (Prakash et al., 2020; Zhao et al., 2024), smart glasses (Lee and Hui, 2018; Xie et al., 2021), and Wi-Fi signals (Ma et al., 2018; Wang et al., 2016a, 2017a). With accurate HAR, systems can anticipate user needs, improve safety, and provide context-aware assistance, fostering a seamless interaction between individuals and their surroundings.
In this paper, we take a step forward to propose a new HIAU task: action summarization. As shown in Fig. 1, users perform continuous actions indoors, which are identified by recognizing each action’s category, start time, and end time to form a sequence of tuples. These tuples can be processed using large language models (LLMs) to generate action summaries, such as insights into physical activity levels and behavioral patterns. With targeted prompts, the system can also detect anomalies like falls or missed medication. We envision several applications for action summarization: (1) Health Monitoring: providing detailed activity logs to support physical rehabilitation, track fitness goals, or detect health issues such as sedentary behavior or irregular movement patterns. (2) Elderly Care: detecting critical events like falls, identifying missed medication times, and offering behavior insights to caregivers for better elderly support. (3) Smart Home Automation: triggering adaptive responses, such as adjusting room temperature based on activity levels or controlling lighting when specific actions are performed. (4) Personal Assistants: Summarizing daily routines to help users with task planning, reminders, and prioritizing activities, thereby enhancing productivity.
The key technology enabling action summarization is temporal action localization (TAL), which focuses on accurately identifying and localizing the start and end times of actions within untrimmed sensory recordings. This step forms the foundation for generating tuples required for further summarization tasks. Over the years, numerous datasets have been developed to support TAL, primarily focusing on video-based tasks. However, these datasets are not well-suited for summarizing indoor daily activities due to several limitations: (1) Task-Specific Datasets: many datasets, such as YouCook2 (Zhou et al., 2018), Breakfast Dataset (Kuehne et al., 2014), and 50Salads (Stein and McKenna, 2013), focus on specific tasks like cooking or food preparation. While these datasets are valuable within their domains, they lack diversity and fail to generalize to other types of everyday indoor activities. (2) Activity Types and Context: datasets like FineAction (Liu et al., 2022), THUMOS14 (Idrees et al., 2017), and ActivityNet (Caba Heilbron et al., 2015) predominantly feature sports or outdoor activities, which do not represent typical indoor scenarios encountered in daily life. (3) Privacy and Applicability: large-scale datasets, such as HACS Segment (Zhao et al., 2019) and ActivityNet (Caba Heilbron et al., 2015), provide diverse activity coverage but are less applicable to indoor settings due to privacy concerns raised by cameras, which limit their use in monitoring indoor activities. In recent years, indoor-focused datasets have started to emerge. For instance: Ego-ADL (Sun et al., 2024): this dataset uses smartphone IMUs, Wi-Fi and audio for TAL but lacks key motion-sensing information from important body positions like the hands and head. This omission makes it challenging to differentiate fine-grained actions, and its reliance on audio raises significant privacy concerns for indoor monitoring applications. WiFiTAD (Liu et al., 2025): while it provides indoor TAL capabilities using Wi-Fi signals, its dataset includes only three participants in a single scene and seven actions, which limits its generalizability.
To address these limitations, we present XRF V2, a novel dataset designed specifically for summarizing indoor daily activities. XRF V2 integrates data from multiple sensing modalities, including Wi-Fi signals, IMU sensors (smartphones, smartwatches, headphones, and smart glasses), and synchronized video recordings, as illustrated in Fig. 1. This dataset was collected with the participation of 16 volunteers, who performed diverse and continuous action sequences across three distinct indoor environments: the dining room, study room, and bedroom, aiming to support a wide range of applications in action summarization. In total, XRF V2 consists of approximately 16 hours and 16 minutes of multimodal data recordings, comprising 853 annotated action sequences that encompass diverse daily routines, including sleeping, cutting fruit, and taking medicine. Compared to WiFiTAD (Liu et al., 2025), XRF V2 provides a more diverse, realistic, and comprehensive benchmark for the action summarization task.
We further propose the XRFMamba neural network to evaluate the XRF V2 dataset and conduct accurate Temporal Action Localization (TAL) for action summarization. XRFMamba is based on the latest advanced achievements in Mamba, which excels at capturing long-term dependencies in temporal sequences. We choose XRFMamba because, unlike action recognition, TAL involves longer input sequences that contain multiple actions, and Mamba has already demonstrated strong capabilities in handling long-term dependencies in fields like NLP (Gu and Dao, 2023) and CV (Zhu et al., 2024). For example, ActionMamba (Chen et al., 2024) has been successfully applied for video-based TAL tasks. In this paper, we are the first to apply the Mamba network to handle Wi-Fi and IMU data for TAL tasks. Compared with existing methods such as ActionFormer (Zhang et al., 2022), TriDet (Shi et al., 2023), ActionMamba (Chen et al., 2024), TemporalMaxer (Tang et al., 2023), UWiFiAction (Wang et al., 2023), and WiFiTAD (Liu et al., 2025), the proposed XRFMamba achieves superior results, e.g., outperforming the latest WiFiTAD by a large margin of 5.49 points in mAP, while using 35% fewer parameters.
Our contributions are threefold:
(1) XRF V2 Dataset: We have collected the XRF V2 dataset, which consists of 16 participants performing continuous indoor actions across three different environments, totaling 30 unique actions. This dataset provides a rich resource for indoor continuous action understanding tasks and systems, offering valuable data for advancing research in this area.
(2) New Task, Method, and Evaluation Metric: We introduce the novel task of action summarization, a new human indoor action understanding task. We also propose a new evaluation metric, Response Meaning Consistency (RMC), to assess the performance of action summarization. We further design a novel method, XRFMamba, to effectively tackle this task. Compared to existing state-of-the-art methods, XRFMamba demonstrates superior performance in action summarization.
(3) Smart Home Intelligent Agents: We demonstrate that continuous sensory data from wearable and Wi-Fi devices, when combined with large language models (LLMs), can enable the development of intelligent agents within smart homes. These agents, such as personal assistants, health assistants, and home assistants, can offer various services, including task reminders, personalized health monitoring, climate control, and more, contributing to the intelligence of smart home systems.
2. Related Work
2.1. Action Understanding with Wi-Fi Signals
Action understanding with Wi-Fi signals has become a key technology in human sensing. The RSSI (Received Signal Strength Indicator) in Wi-Fi signals is commonly used for indoor localization (Bahl and Padmanabhan, 2000; Wu et al., 2012b). However, its accuracy is relatively low and is susceptible to multipath effects and temporal dynamics (Yang et al., 2013; Wu et al., 2011). As a result, later works have adopted more information-rich Channel State Information (CSI) to achieve more precise localization (Yang et al., 2013), such as FILA (Wu et al., 2012a), SpotFi (Kotaru et al., 2015), Splicer (Xie et al., 2015), and Dynamic-MUSIC (Li et al., 2016a). In addition to high-precision indoor localization, Wi-Fi CSI signals have been widely applied in various tasks in fields such as smart homes, health monitoring, and human-computer interaction. These include hand gesture recognition (Pu et al., 2013; Ma et al., 2018), finger gesture recognition (Li et al., 2016b; Tan and Yang, 2016; Ali et al., 2015), activity recognition (Wang et al., 2015, 2016a, 2014), vital sign monitoring (Wang et al., 2016b, 2017b; Yu et al., 2021), pose estimation (Jiang et al., 2020; Wang et al., 2019b; Yan et al., 2024; Wang et al., 2019a), mesh reconstruction (Wang et al., 2022b, c), crowd counting (Xi et al., 2014), and human imaging (Xu et al., 2024; Wang et al., 2019b). Since Wi-Fi-based human sensing is highly sensitive to environmental changes, several studies have concentrated on enhancing its cross-environment capabilities. For example, EI (Jiang et al., 2018) and Person-in-WiFi (Wang et al., 2019b) use domain adaptation techniques to align the features of the source and target environments for activity recognition and pose estimation, respectively. CrossSense (Zhang et al., 2018) proposes using a mixture of experts approach to train multiple tasks and select the optimal expert for the target environment, maintaining cross-environment adaptability. Widar 3.0 (Zheng et al., 2019), on the other hand, proposes environment-invariant representation, i.e., body velocity profiles, to achieve cross-environment, cross-person, and cross-method hand gesture recognition.
Apart from the tasks mentioned above, some researchers have also focused on understanding continuous action using Wi-Fi. Smokey (Zheng et al., 2016) recognizes smoking actions by detecting periodic patterns, such as holding the cigarette, bringing it to the mouth, and inhaling the smoke. WiLife (Li et al., 2024a) uses Wi-Fi for continuous in-home monitoring of elderly individuals living alone, ensuring their safety and well-being. Wang et al. (Wang et al., 2023) employ U-shaped deep neural networks for the temporal localization and segmentation of daily activities and human-human interactions. WiFiTAD (Liu et al., 2025) introduces a frequency-aware learning approach combined with a dual-pyramid network to promote the accuracy of temporal action localization. IMar (He and Yang, 2022) employs tensor decomposition to extract individual Wi-Fi streams for each person, enabling separate recognition of continuous actions. In comparison to these works, we push the boundaries further by introducing a new task: action summarization. We envision this task supporting applications in remote healthcare and monitoring, personal assistants, smart home automation, and more.
2.2. Action Understanding with IMUs
Inertial Measurement Units (IMUs), such as accelerometers and gyroscopes, can measure motion acceleration and angular velocity. They are widely embedded in personal consumer electronics, including smartphones, smartwatches, smart glasses, and earbuds. Researchers have leveraged IMUs in these devices to develop numerous applications in areas such as healthcare (Alam, 2020; Jung et al., 2022; DeVrio et al., 2023; Shang et al., 2024; Wang et al., 2024b; De Brabandere et al., 2020; Wang et al., 2022d; Meng et al., 2022), sports (Tian et al., 2021; Wolff et al., 2018), human-computer interaction (Laput and Harrison, 2019; DeVrio et al., 2023; Sun et al., 2024; DeVrio et al., 2023; Wahl et al., 2015; Scholl et al., 2015; Verma et al., 2021; Zhao et al., 2024), and industrial applications (Azadi et al., 2019; Sopidis et al., 2022). Jung et al. (Jung et al., 2022) utilize smartwatches and continuous wavelet transform (CWT) technology to identify tremor episodes and amplitude of Parkinson’s disease patients during drawing tasks. Azadi et al. (Azadi et al., 2019) investigate the feasibility of recognizing screwing actions in manufacturing using smartwatches and unsupervised learning techniques. De et al. (De Brabandere et al., 2020) place a mobile phone on a patient’s left hip to estimate hip and knee joint loading, providing valuable insights for doctors in diagnosing the patient’s condition. Anguita et al. (Anguita et al., 2013) leverage a smartphone worn at the waist to identify six daily activities: standing, sitting, lying down, walking, walking downstairs, and walking upstairs. WISEGlass (Wahl et al., 2015) captures the wearer’s head movements—up, right, left, and down—using IMUs embedded in the legs of smart glasses, enabling the control of the Pac-Man game accordingly. Wetlab (Scholl et al., 2015) integrates Google Glass with a wrist-worn IMU sensor, allowing for the capture and review of DNA extraction experiments conducted in wet laboratory environments. ExpressEar (Verma et al., 2021) uses earbuds to detect fine facial muscle movements, exploring the correlation between movements and facial expressions. This enables hands-free control through intentional facial expressions or monitoring involuntary ones to infer the user’s emotional state. ExpressEar (Verma et al., 2021) utilizes a Nokia eSense earbuds to sense fine facial muscle movements, and mine the correlation between the movements and facial expressions for hands-free control using facial expressions or to monitor involuntary facial expressions to infer the user’s emotional state. Ui-Ear (Zhao et al., 2024) utilizes earbuds to detect on-face gestures, including tapping, slide up, slide down, slide left, slide right, clockwise circle, and counterclockwise circle, enhancing the user-earbuds interaction experience.
Current advancements primarily focus on Human Activity Recognition (HAR), but a recent work (Bock et al., 2024) shifts its focus to Temporal Action Localization (TAL) using IMU data. It adapts video-based TAL methods, e.g., ActionFormer (Zhang et al., 2022), TemporalMaxer (Tang et al., 2023), and TriDet (Shi et al., 2023) on multiple public datasets, including WetLab for DNA extraction experiment (Scholl et al., 2015) and Hang-Time HAR for basketball-playing actions (Hoelzemann et al., 2023). We take a step further by collecting a continuous dataset of indoor daily activities, specifically designed to evaluate both Temporal Action Localization (TAL) and action summarization tasks.
2.3. Continuous Action Dataset and Benchmark
Temporal action localization (TAL) and temporal action detection (TAD) are fundamental tasks in the computer vision community and have been extensively studied (Shou et al., 2016). Searchers have introduced numerous publicly accessible datasets to evaluate TAL and TAD methods. Breakfast (Kuehne et al., 2014) and YouCook2 (Zhou et al., 2018) are two representative large-scale daily cooking datasets captured from third-person viewpoint cameras. The Breakfast includes 10 operations related to breakfast preparation, performed by 52 different people from 18 different kitchens. YouCook2 contains 2000 untrimmed long videos from 89 cooking recipes, with each step annotated with temporal boundaries and described in sentences. Beyond these cooking datasets, some other datasets such as ActivityNet (Caba Heilbron et al., 2015), THUMOS14 (Idrees et al., 2017), HACS Segments (Zhao et al., 2019), FineAction (Liu et al., 2022), are released for TAL involving daily activities, sports, and household tasks. Among these, FineAction consists of 103K temporal instances of 106 action categories, annotated in 17K untrimmed videos, covering household activities, personal care, socializing, and sports. While cameras are widely used for TAL tasks, they face challenges such as fixed positions and viewing angles once deployed, as well as privacy concerns in home environments. Thus, several studies have explored smartwatches, smartphones, and Wi-Fi for TAL. UWash (Wang et al., 2024b) is a dedicated system for segmenting handwashing gestures using smartwatch IMU data. Its dataset includes 550 handwashing recordings from 71 participants. Ego-ADL (Sun et al., 2024) leverages a smartphone to continuously capture users’ audio, Wi-Fi, and motion sensor data in daily life. The released dataset comprises 141 annotated sequences. However, the reliance on a single smartphone placed in the user’s pocket introduces limitations in capturing fine-grained hand and head movements. WiFiTAD (Liu et al., 2025) uses a setup with one transmitter and one receiver Wi-Fi device to collect data in a single indoor environment. The dataset features recordings from three participants performing seven daily activities—walking, running, jumping, waving, falling, sitting, and standing. It includes 553 untrimmed sequences, with an average of 3.79 action instances per sequence.
Compared to UWash, Ego-ADL, and WiFiTAD, XRF V2 offers richer sensor data, integrating two smartphones, two smartwatches, one pair of earbuds, and one smart glasses. This comprehensive setup captures diverse motion information, including body, hand, and head movements, enabling more effective Temporal Action Localization (TAL) in home environments. Since WiFiTAD is the most relevant and latest work to XRF V2, we compare XRF V2 with it in Table 1, which demonstrates significant advantages over WiFiTAD in terms of the number of scenes, actions, subjects, sequences, and modalities. Furthermore, XRF V2 supports the exploration of optimal sensor combinations and multimodal data fusion, providing a comprehensive dataset for advancing IMU and Wi-Fi based temporal action localization and action summarization research.
Table 1. XRF V2 shows great advantages in the number of scenes, actions, subjects, sequences, and modalities compared to the very latest WiFiTAD (Liu et al., 2025).
Dataset #Scene #Actions #Subjects #Sequences Modality Venue
WiFiTAD (Liu et al., 2025) 1 7 3 553 Wi-Fi AAAI 2025
XRF V2 3 30 15 853 (16h 16m 8s) Wi-Fi, 5 IMUs, RGB+D+IR /
3. XRF V2 Acquisition and Description
3.1. Hardware Setups
Wi-Fi transceivers. The Wi-Fi transceiver devices include one transmitter and three receivers, positioned at the four corners of a 3.1 m $\times$ 2.4 m rectangular area. This rectangle defines the activity region for human actions, and the devices are placed at a height of 120 cm to ensure effective coverage for full-body motion capture. The transmitter operates with a single antenna, broadcasting packets at 200 per second on channel 128 (5.64 GHz). Each receiver is equipped with three antennas. Using the Wi-Fi CSI Tool (Halperin et al., 2011), we extract the channel state information (CSI) from 30 orthogonal frequency-division multiplexing (OFDM) subcarriers. The resulting CSI recordings are represented as a tensor of size $(200t)\times 1\times 3\times 3\times 30$, where $t$ is the recording duration in seconds. For example, for an 80-second continuous action, the CSI recordings yield a tensor of size $16000\times 1\times 3\times 3\times 30$, capturing detailed variations in the wireless channel over time for subsequent analysis.
IMU devices. The IMU device choice follows IMUPoser (Mollyn et al., 2023) and includes three commonly used devices in daily life: smartwatches, smartphones, and earbuds. Additionally, we introduce smart glasses, considering their promising potential in AR and the metaverse (ZHAO et al., 2023). In our experiments, the brand of earbuds is AirPods Pro, with IMU data recorded at a sampling rate of 25 Hz. For smartwatches, smartphones, and smart glasses, we use commercially available IMU modules to simulate the IMUs of these devices. These modules allow centralized control through a PC-based interface, facilitating better synchronization across all IMU devices. The IMU modules record data at a sampling rate of 50 Hz. The placement of IMU devices is shown in Fig. 1.
Azure Kinect. The Azure Kinect captures RGB, depth, and infrared images at 30 frames per second in 720P resolution using its toolkit. The Kinect is placed between the transmitter and the third receiver at a height of 115 cm to capture video footage of the movement of individuals in the scene. We can use open-source computer vision models to extract pose and mesh information from the videos, which will be useful for future tasks such as IMU- and Wi-Fi-based pose estimation and mesh reconstruction.
The experimental scenarios are shown in Fig. 2. Before data recording, all the aforementioned hardware devices undergo careful synchronization. Notably, XRF V2 selects commonly available consumer-grade products used in everyday household and personal settings, unlike XRF55 (Wang et al., 2024a), which utilizes millimeter-wave radar and an RFID reader with a tag array.
Figure 2. Illustration of the experimental scenarios, where volunteers perform continuous action sequences in the dining room, study room, and bedroom. IMU devices from two smartwatches, two smartphones, a pair of earbuds, a pair of glasses, and four Wi-Fi transceivers, along with an Azure Kinect, are synchronized to record action sequences.
3.2. Action Sequence Proposal
We select three typical household scenarios for the experiment: the dining room, study room, and bedroom. These settings allow us to summarize users’ activities related to eating, studying, and sleeping, along with associated series of actions. For example, activities such as taking medicine, drinking water, washing hands, and stretching serve as health-related indicators; and actions like opening curtains or turning on lights provide insights into interactions with household devices.
We communicate the above objective to the volunteers, but instead of assigning each person a fixed action sequence, we aim to increase the diversity of action sequences and incorporate personal habits relevant to the volunteers. Therefore, we ask each volunteer to recall their own daily action sequences in the dining room, study room, and bedroom, and estimate the execution time of each action in the sequence. Each volunteer recalls five action sequences based on the starting positions of the actions, with each sequence containing 7-10 actions. Fig. 4 shows examples of some action sequences, while Table 2 presents the proposed actions in each scene.
Table 2. Actions proposed and executed by volunteers in the Bedroom, Study room, and Dining room.
Scene Proposed actions
Dining room 1. Walk (towards the cabinet, towards the chair), 2. Sit down, 3. Stand up, 4. Pour water into the cup, 5. Drink water, 6. Take medicine, 7. Pick up things, 8. Take the fruits from the cabinet, 9. Cut fruits, 10. Eat fruits, 11. Wash hands, 12. Throw waste, 13. Wipe the table, 14. Stretch when sitting.
Study room 1. Walk (towards the door, towards the chair, towards the blackboard), 2. Sit down, 3. Stand up, 4. Pour water into the cup, 5. Drink water, 6. Take medicine, 7. Turn on and off the desk lamp, 8. Operate the mouse, 9. Write, 10. Operate the keyboard, 11. Read a book, 12. Open an envelope, 13. Answer the phone, 14. Stretch when standing, 15. Write on the blackboard.
Bedroom 1. Walk (towards the bed, towards the window), 2. Get up, 3. Sit down, 4. Lie down, 5. Stand up, 6. Read a book, 7. Pour water into the cup, 8. Drink water, 9. Take medicine, 10. Stretch when standing, 11. Use phone, 12. Open and close windows, 13. Open and close curtains, 14. Water plants, 15. Stand still, 16. Lying still.
3.3. Volunteer Coordination and Action Execution
We recruit volunteers through group chats, online meetings, and social media. Volunteers can access the informed consent form online, which includes details about the study’s purpose, methods, participation duration, potential risks, and compensation. Registrants receive a shared editable online spreadsheet to fill in their preferred participation time. At the experiment site, volunteers are provided with a printed version of the informed consent form. We explain its content to each volunteer and address questions they might have. If volunteers agree to proceed, they sign to confirm their participation and receive compensation equivalent to the local average wage multiplied by their participation duration.
Before starting the experiment, we assist volunteers in wearing the IMU devices and explain the data collection process and relevant precautions. Once the volunteers feel ready, the experiment begins. Volunteers still retain the right to withdraw from the experiment at any time. Our experiment is conducted with approval from the Institutional Review Board (IRB). We recruit 16 volunteers, with an age range of 22–34 years, a height range of 1.57–1.82 meters, and a weight range of 43–90 kgs.
Before performing the action sequences, we proportionally shorten the durations estimated by the volunteers to fall within 5–20 seconds. This adjustment is necessary because the originally estimated durations are often much longer; for instance, reading a book might be estimated as one hour. Performing such lengthy actions would significantly increase data collection time and complicate annotation. For example, it is challenging for individuals to maintain a single activity, such as reading, for an extended period without engaging in other incidental actions, such as adjusting their posture or picking up a drink. When these additional actions occur, the annotation process becomes more complicated, as it requires meticulously reviewing the entire video segment to precisely identify and mark the start and end times of each action.
Figure 3. Automatic annotation process. The audio prompts are played in real-time to guide the volunteers. Upon hearing an action instruction, the volunteer performs the corresponding action and transitions to the next action upon hearing the subsequent prompt. The start time and end time of an action are annotated with the prompts played time.
We also make reasonable adjustments to the execution durations several times, generating more action sequences. As a result, the action sequences in each scene become more diverse. After that, we can adopt an automated annotation method to streamline the process. The action sequences proposed by the volunteers are converted into audio prompts using text-to-audio technology. These audio prompts are played in real time to guide the volunteers. Upon hearing an action instruction, the volunteer performs the corresponding action and transitions to the next action upon hearing the subsequent prompt. This process continues until the entire sequence is completed.
The start and end times of each audio prompt are automatically recorded as the start and end times of the corresponding action. Fig. 3 illustrates this automatic annotation process. As a result, after executing an action sequence containing $n$ actions, we automatically obtain the annotations for the entire sequence: $\{(a_{i},t_{i}^{s},t_{i}^{e}),i=1,2,\dots,n\}$, where $a_{i}$, $t_{i}^{s}$, and $t_{i}^{e}$ represent the action category, start time, and end time of the $i$-th action, respectively.
In general, the duration of a single action sequence is approximately 50–80 seconds. After completing one action sequence, volunteers rest for 10 seconds, and after executing five sequences, they take a 5-minute break.
Figure 4. Examples of action sequence from the study room, dining room, and bedroom, with a detailed visualization of the sensory sequence from the bedroom.
In Fig. 4, we present action sequences from the study room, dining room, and bedroom, with a sensory sequences visualization from the bedroom. For clarity, we display only the three-axis accelerometer data from all IMU devices, and for the Wi-Fi data, we select the data from one antenna of each receiving device. Additionally, the sampling frequency of all data has been standardized, and the x-axis represents the timestamp. To highlight the different actions in the sequence, we use dashed lines in the figure to indicate the endpoints of each action.
3.4. Dataset Statistics
Each of the 16 volunteers proposed 5 action sequences for each scene. By shortening durations and adjusting the length of individual actions, we generated multiple variants of these sequences. Specifically, in the bedroom, each volunteer performed three different variations of each action sequence, while in the study room and dining room, four different variations were performed for each sequence. Thus, XRF V2 dataset contains a total of: $16\times 5\times 3+16\times 5\times 4+16\times 5\times 4=880$ action sequences. After filtering out incomplete and erroneous sequences, we obtained 853 valid action sequences, with a total duration of 16 hours, 16 minutes and 8 seconds. These sequences were captured using synchronized Channel State Information (CSI) from four Wi-Fi transceivers, IMU data from two smartwatches, two smartphones, a pair of earbuds, and a pair of smart glasses, along with RGB+D+IR video streams from a Kinect camera. Furthermore, we implemented a stratified partitioning strategy maintaining an 80:20 ratio between training and testing sets across all environmental contexts and participants, resulting in 682 sequences (80%) allocated for model training and 171 sequences (20%) reserved for evaluation purposes. Table. 3 shows detailed statistics of XRF V2 dataset.
Table 3. XRF V2 contains 853 multimodal sensory sequences for action summarization and temporal action localization.
train test
Dining room 244 sequences 4h39min41s 61 sequences 1h8min40s
Study room 256 sequences 5h0min31s 64 sequences 1h16min15s
Bedroom 182 sequences 3h19min8s 46 sequences 51min53s
Total 682 sequences 12h59min20s 171 sequences 3h16min48s
4. Method
State Space Models (SSMs), such as S4 (Gu et al., 2021), S5 (Smith et al., 2022), and Mamba (Gu and Dao, 2023), are a recent class of deep learning methods that have demonstrated advantages in both natural language processing (Gu and Dao, 2023) and computer vision (Zhu et al., 2024). Unlike Transformer-based models, which suffer from quadratic time complexity due to self-attention, SSMs operate with linear-time complexity, making them significantly more efficient for long sequence modeling. In action summarization, the input sensor sequences are much longer than those used in action recognition tasks, necessitating the ability to effectively model long-range dependencies. To this end, we base our approach on an advanced SSM model and propose the XRFMamba network. Next, we first introduce the preliminary concepts of SSMs in Sec. 4.1, then, we will present XRFMamba in Sec. 4.2.
4.1. State Space Model Preliminaries
Figure 5. State space model (a). The model can be computed in linear recurrence (b) for inference, or global convolution (c) for training.
As introduced in Mamba (Gu and Dao, 2023), state space models are inspired by a continuous system, that maps 1-dimensional function or sequence $x(t)\in\mathbb{R}\mapsto y(t)\in\mathbb{R}$ through a hidden state $h(t)\in\mathbb{R}^{\mathtt{N}}$, illustrated in Fig. 5 (a). The system works as follows:
(1) $h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),\hskip 10.0pty(t)=\mathbf{C}h(t).$
where $\mathbf{A}\in\mathbb{R}^{\mathtt{N}\times\mathtt{N}}$ is the evolution parameter, and $\mathbf{B}\in\mathbb{R}^{\mathtt{N}\times 1}$ and $\mathbf{C}\in\mathbb{R}^{1\times\mathtt{N}}$ are the projection parameters.
For practical implementation, the continuous system needs to be discretized. When discretizing, the continuous state transition matrices $(\mathbf{A},\mathbf{B})$ are approximated by their discrete counterparts $(\mathbf{\overline{A}}$, $\mathbf{\overline{B}})$. The Zero-Order Hold (ZOH) is commonly used in the discretization process, defined as follows:
(2) $\mathbf{\overline{A}}=\exp{(\Delta\mathbf{A})},\hskip 10.0pt\mathbf{\overline{B}}=(\Delta\mathbf{A})^{-1}(\exp{(\Delta\mathbf{A})}-\mathbf{I})\cdot\Delta\mathbf{B}.$
After the discretization, Equation (1) can be rewritten as:
(3) $h_{t}=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t},\hskip 10.0pty_{t}=\mathbf{C}h_{t}.$
Since this is in the linear recurrence expression, as illustrated in Fig. 5 (b), it is efficient for autoregressive inference.
In addition, the discretized SSM can also be expressed in the global convolution expression, as follows:
(4) $\mathbf{\overline{K}}=(\mathbf{C}\mathbf{\overline{B}},\mathbf{C}\mathbf{\overline{A}}\mathbf{\overline{B}},\dots,\mathbf{C}\mathbf{\overline{A}}^{\mathtt{L}-1}\mathbf{\overline{B}}),\hskip 10.0pt\mathbf{y}=\mathbf{x}*\mathbf{\overline{K}},$
where $\mathtt{L}$ is input sequence length, and $\overline{\mathbf{K}}\in\mathbb{R}^{\mathtt{L}}$ is a structured convolutional kernel, illustrated in Fig. 5 (c). The convolution expression is efficient for parallel training.
Due to the flexibility of SSMs, they can switch to convolution during training and to recurrence during inference, providing fast performance in both stages. SSMs have a rich foundational knowledge, and introducing all aspects is beyond the scope of this paper. For more detailed information, please refer to Mamba (Gu and Dao, 2023). In practice, SSMs are highly convenient. For example, in Vision Mamba (Zhu et al., 2024), the Mamba module replaces the Transformer module in the Vision Transformer (ViT) (Dosovitskiy et al., 2021), achieving better performance than ViT while also being more computationally efficient. Additionally, in the Video Mamba Suite (Chen et al., 2024), Mamba is evaluated as a plug-in or backbone, showing excellent results across 12 video understanding tasks, including action recognition and temporal action localization.
4.2. XRFMamba
To achieve action summarization, it is essential to identify the actions performed by the user within a specific time span, along with their respective start and end times, which corresponds to the Temporal Action Localization (TAL) task. To address TAL, we propose XRFMamba, a novel neural network based on the State Space Model (SSM). As shown in Fig. 6, XRFMamba employs Mamba blocks as the representation learning backbone, enabling the accurate localization of actions in continuous sensory sequences.
Figure 6. XRFMamba uses Mamba as the feature learning backbone. It takes an input sequence and predicts the action categories that appear in the sequence, along with the start and end times of each action.
Projection and Fusion. The Projection module is adapted from WiFiTAD (Liu et al., 2025) and comprises three sequential components: 1D Convolutional Operations, Group Normalization, and ReLU Activation. The Projection module standardizes the output channels of IMU and Wi-Fi data using convolutional operations, ensuring consistent channel dimensions for the subsequent fusion operation. The fusion of Wi-Fi and IMU data is performed through a weighted summation, as shown in the following equation:
(5) $e=\lambda e_{wifi}+(1-\lambda)e_{imu}$
Here, $e_{wifi}$ and $e_{imu}$ represent Wi-Fi and IMU data outputs after projection, respectively. In our experiments, we observed that the IMU data performed better than Wi-Fi data. To maintain the stability of the fusion performance, we assigned a higher weight to the IMU data. Specifically, we set $\lambda=0.2$ in our experiments.
Embedding. After fusion, the design of our embedding module takes inspiration from ActionMamba (Chen et al., 2024). In ActionMamba, the input to the Mamba backbone is not raw video but feature embeddings derived from the pre-trained model, InterVideo (Wang et al., 2022a). Following a similar approach, we introduce an embedding step prior to feeding the data into the Mamba backbone. However, due to the absence of widely accepted pre-trained models for IMU and Wi-Fi data in both academic and industrial contexts, a dedicated module is required to generate these embeddings. In this regard, we leverage the TSSE (Temporal Signal Semantic Embedding) module from WiFiTAD, which has proven effective in embedding Wi-Fi data for temporal action localization tasks. The TSSE embedding module is a dual-stream architecture, where one stream is based on the Transformer, which excels at learning global features, and the other is based on convolutional layers, which are effective at capturing local features. These two streams are fused at the output stage to create a comprehensive feature representation. The fused embeddings are then passed into the Mamba backbone for further processing.
Mamba Backbone. The Mamba backbone in Fig.6 is the Decomposed Bidirectionally Mamba (DBM), introduced in ActionMamba (Chen et al., 2024). DBM extends the original Vision Mamba (ViM) (Zhu et al., 2024) and vanilla Mamba (Gu and Dao, 2023). We present DBM block details in Algorithm. 31, consistent with the ViM algorithm style. Initially, the input sensory sampling point $\mathbf{S}_{l-1}$ is normalized and processed through four linear layers to produce $\mathbf{z_{1}}$, $\mathbf{x_{forward}}$, $\mathbf{x_{backward}}$, and $\mathbf{z_{2}}$. $\mathbf{x_{forward}}$ and $\mathbf{x_{backward}}$ are then processed in their respective forward and backward directions. Specifically, $\mathbf{x_{forward}}$ undergoes forward processing to yield $\mathbf{x^{\prime}}_{forward}$, $\mathbf{B}_{forward}$, and $\mathbf{C}_{forward}$, which are subsequently discretized into $\overline{\mathbf{A}}_{forward}$ and $\overline{\mathbf{B}}_{forward}$. Next, $\mathbf{y_{forward}}$ is computed through the forward SSM and fused with $\mathbf{z_{1}}$ to generate $\mathbf{y_{forward}^{{}^{\prime}}}$. Similarly, $\mathbf{y_{backward}^{{}^{\prime}}}$ is obtained in the backward direction. The forward and backward outputs, $\mathbf{y_{forward}^{{}^{\prime}}}$ and $\mathbf{y_{backward}^{{}^{\prime}}}$, are then fused to produce the final output $\mathbf{S}_{l}$. Notably, the convolution and SSM parameters are shared between both directions in DBM.
Algorithm 1 Decomposed Bidirectionally Mamba Block Process (in ViM (Zhu et al., 2024) style)
0:  sensory sampling point $\mathbf{S}_{l-1}$ : $(\mathtt{B},\mathtt{L},\mathtt{D})$
1:  /* normalize the input sequence $\mathbf{S}_{l-1}^{\prime}$/
2:  $\mathbf{S}_{l-1}^{\prime}$ : $(\mathtt{B},\mathtt{L},\mathtt{D})$leftarrow $\mathbf{LayerNorm}(\mathbf{S}_{l-1})$
3:  $\mathbf{z1}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{1}}(\mathbf{S}_{l-1}^{\prime})$
4:  $\mathbf{x_{forward}}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{2}}(\mathbf{S}_{l-1}^{\prime})$
5:  $\mathbf{x_{backward}}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{3}}(\mathbf{S}_{l-1}^{\prime})$
6:  $\mathbf{z2}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow $\mathbf{Linear}^{\mathbf{4}}(\mathbf{S}_{l-1}^{\prime})$
7:  /* process with different direction */
8:  for $o$ in {forward, backward} do
9:     $\mathbf{x}^{\prime}_{o}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow $\mathbf{SiLU}(\mathbf{Conv1d}_{o}(\mathbf{x_{o}}))$
10:     $\mathbf{B}_{o}$ : $(\mathtt{B},\mathtt{L},\mathtt{N})$leftarrow $\mathbf{Linear}_{o}^{\mathbf{B}}(\mathbf{x}^{\prime}_{o})$
11:     $\mathbf{C}_{o}$ : $(\mathtt{B},\mathtt{L},\mathtt{N})$leftarrow $\mathbf{Linear}^{\mathbf{C}}_{o}(\mathbf{x}^{\prime}_{o})$
12:     /* softplus ensures positive $\Delta_{o}$/
13:     $\Delta_{o}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow $\log(1+\exp(\mathbf{Linear}_{o}^{\Delta}(\mathbf{x}^{\prime}_{o})+\mathbf{Parameter}_{o}^{\Delta}))$
14:     /* shape of $\mathbf{Parameter}_{o}^{\mathbf{A}}$ is $(\mathtt{E},\mathtt{N})$/
15:     $\overline{\mathbf{A}_{o}}$ : $(\mathtt{B},\mathtt{L},\mathtt{E},\mathtt{N})$leftarrow $\Delta_{o}\bigotimes\mathbf{Parameter}_{o}^{\mathbf{A}}$
16:     $\overline{\mathbf{B}_{o}}$ : $(\mathtt{B},\mathtt{L},\mathtt{E},\mathtt{N})$leftarrow $\Delta_{o}\bigotimes\mathbf{B}_{o}$
17:     /* initialize $h_{o}$ and $\mathbf{y}_{o}$ with 00 */
18:     $h_{o}$ : $(\mathtt{B},\mathtt{E},\mathtt{N})$leftarrow zeros $(\mathtt{B},\mathtt{E},\mathtt{N})$
19:     $\mathbf{y}_{o}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow zeros $(\mathtt{B},\mathtt{L},\mathtt{E})$
20:     /* SSM recurrent */
21:     for $i$ in {0, …, M-1} do
22:        $h_{o}$ = $\overline{\mathbf{A}_{o}}[:,i,:,:]\bigodot h_{o}+\overline{\mathbf{B}_{o}}[:,i,:,:]\bigodot\mathbf{x}_{o}^{\prime}[:,i,:,{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\mathtt{None}}]$
23:        $\mathbf{y}_{o}[:,i,:]$ = $h_{o}\bigotimes\mathbf{C}_{o}[:,i,:]$
24:     end for
25:  end for
26:  /* get gated $\mathbf{y}$/
27:  $\mathbf{y_{forward}^{{}^{\prime}}}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow $\mathbf{y_{forward}}\bigotimes\mathbf{SiLU}(\mathbf{z1})$
28:  $\mathbf{y_{backward}^{{}^{\prime}}}$ : $(\mathtt{B},\mathtt{L},\mathtt{E})$leftarrow $\mathbf{y_{backward}}\bigotimes\mathbf{SiLU}(\mathbf{z2})$
29:  /* residual connection */
30:  $\mathbf{S}_{l}$ : $(\mathtt{B},\mathtt{L},\mathtt{D})$leftarrow $\mathbf{Linear}^{\mathbf{S}}(\mathbf{Cat}(\mathbf{y_{forward}^{{}^{\prime}}},\mathbf{y_{backward}^{{}^{\prime}}}))+\mathbf{S}_{l-1}$
31:  Return: $\mathbf{S}_{l}$ : $(\mathtt{B},\mathtt{L},\mathtt{D})$
Pyramidal Prediction Head. We use Mamba to transform the input into features with the same number of channels but different lengths, resulting in a feature pyramid composed of $(f_{1},f_{2},\dots,f_{K})$. Each feature dimension $f_{i}$ is processed through two separate convolutional paths to output action recognition results and action localization results, respectively. The total loss is the sum of the errors from these two prediction branches. The overall training loss of the network is given as follows:
(6) $L^{i}$ $=\alpha_{1}L_{Cls}^{i}+\alpha_{2}L_{Loc}^{i}$
$L$ $=\sum_{i=1}^{K}L^{i}$
where we use Focal Loss (Lin et al., 2017b) for action classification and $L_{1}$ loss for action localization; $\alpha_{1}$ and $\alpha_{2}$ are used to balance the scale of the two losses, and we set $\alpha_{1}=1$ and $\alpha_{2}=1000$ based on empirical observations to ensure that both losses are on a similar magnitude. This pyramidal prediction strategy leverages the ability of different lengths to capture actions of varying durations, originally applied in object detection (Lin et al., 2017a), where features of different sizes capture objects of different scales, and has been widely adopted in temporal action localization tasks, such as in (Zhang et al., 2022; Liu et al., 2025; Chen et al., 2024).
Figure 7. Each action has multiple localization priors, leading to several candidate prediction results. NMS enables XRFMamba to keep only the best result for each action. (Illustration inspired by RF-fall (Tian et al., 2018))
Inference. In XRFMamba, each action has multiple localization priors, leading to several candidate prediction results. To convert multiple localization prediction results into a single prediction, we use Non-Maximum Suppression (NMS). The steps of NMS are as follows: (1) Sorting: sort all candidate segments by their scores, processing the segment with the highest score first. (2) Discard Overlapping Segments: after processing the current segment, compute its overlap with other segments. If the overlap exceeds a predefined threshold, that segment is discarded. (3) Repeat the Process: continue with the next highest-scoring segment and repeat step 3 until all segments have been processed. Fig. 7 shows an example. The segment corresponding to the highest classification score for action $i$ is selected first, and the other segments that have a high overlap with it are discarded. As a result, action $i$ will only retain one localization prediction. Then, among the remaining segments, the segment with the highest score for action $j$ is selected, and all the other segments with a high overlap with it are discarded. Therefore, action $j$ will also have only one prediction result. NMS enables XRFMamba to keep only the best result for each action.
4.3. Implementation Details
Prior to training, we perform data augmentation on action sequence sensory data in the training set through sliding window segmentation. Each 30-second clip is generated with a 3-second sliding stride, expanding the original 682 training action sequences into 9660 30-second clips. Subsequently, we unify multimodal sensor clips to 2048 temporal dimensions via signal interpolation: (1) Earbuds clips (25Hz×$\times$30s = 750 samples) and other IMU clips (50Hz×$\times$30s = 1500 samples) are upsampled to 2048 using linear interpolation; (2) Wi-Fi clips (200Hz×$\times$30s = 6000 samples) are adaptively downsampled to 2048.
The sliding-window data augmentation introduces inherent challenges of partial action truncation, as 30-second clips contain multiple action instances. Temporal boundaries of clips frequently manifest partial action executions (head/tail truncation), where left boundaries preserve end time without start time, while right boundaries retain start time lacking end time. To address this truncation issue, we (1) incomplete segments covering $\geq 80\%$ of original action duration are retained, with truncation points redefined as start/end time; (2) Segments containing $<80\%$ complete actions are excluded from loss computation via temporal masking during backpropagation.
We use AdamW optimizer to train XRFMamba for 80 epochs on one Nvidia 3090 GPU. The initial learning rate is 4e-5, and is adjusted every 30 epochs by a factor of 0.5. The batch size is 8.
During the inference phase, XRFMamba processes 30-second temporal clips with a 5-second sliding stride, generating preliminary action proposals for each clip. To consolidate overlapping predictions, we implement the temporal non-maximum suppression (NMS) algorithm described above. The final temporal action localization (TAL) outputs comprise action classifications with precise start/end time boundaries.
5. Evaluation and Results
5.1. Metrics
(1) AP and mAP for Temporal Action Localization. We use the intersection over union (IoU) between the predicted action time and the ground truth time, referred to as tIoU, to evaluate the performance of Temporal Action Localization (TAL). A higher tIoU indicates better performance. tIoU can be computed as follows:
(7) $tIoU=\frac{time~{}of~{}overlap}{time~{}of~{}uniou}=\frac{\mathtt{max(0,\mathtt{min}(end_{pred},end_{gt}))}-\mathtt{max(0,\mathtt{min}(start_{pred},start_{gt}))}}{\mathtt{end_{pred}-start_{pred}+end_{gt}-start_{gt}}}$
where $\mathtt{start_{pred}}$ and $\mathtt{end_{pred}}$ are the predicted start and end times of an action; $\mathtt{start_{gt}}$ and $\mathtt{end_{gt}}$ are the ground truth start and end times.
Furthermore, we use AP@t to denote the ratio of actions whose tIoU is greater than or equal to a threshold $t$, computed as follows:
(8) $AP@t=\frac{\sum_{i=1}^{N}\mathtt{I}(tIoU_{i}\geq t)}{N}$
where $N$ is the number of actions to compute tIoU. $\mathtt{I}$ outputs 1 if $IoU_{i}\geq t$ is true, else 0. The larger the AP@t, the higher the ratio of actions with a tIoU greater than or equal to the threshold $t$. On the other hand, as $t$ increases, the threshold for achieving a higher tIoU becomes more challenging, leading to a decrease in AP@t. The upper bound of AP@t is 1, while the lower bound is 0.
Besides, we use mAP (mean Average Precision) as the ultimate metric to evaluate the accuracy of Temporal Action Localization (TAL). It is computed by averaging the AP@t over 10 tIoU levels, specifically AP@[0.50:0.05:0.95], which covers tIoU thresholds from 0.5 to 0.95 with a step size of 0.05.
Figure 8. Example of Response Meaning Consistency (RMC) calculation. For each test action sequence, several actions with start and end times are provided for both ground truth and predicted values. Using a task-oriented prompt, the LLM generates corresponding action summarization results. Human auditors then assess whether the responses for the real and predicted values convey consistent meanings. A higher level of consistency results in a higher RMC score.
(2) RMC and mRMC for Action Summarization. Action summarization involves providing a large language model (LLM) with a sequence of actions along with their start and end times. The task is then to use task-oriented prompts to generate the responses. This is a novel task, and there is currently no standard metric for evaluating the effectiveness of action summarization. To address this, we introduce a new metric, Response Meaning Consistency (RMC), to evaluate action summarization performance.
As shown in Fig. 8, an action sequence (as) contains both the ground truth actions and their corresponding start and end times. XRFMamba also predicts action categories along with their start and end times. After inputting the sequence into the LLM for question-answering, the LLM generates responses based on both the ground truth and the predictions. When the model’s predictions are more accurate, the responses generated for the ground truth and the predictions should be consistent in meaning. To evaluate this, we invite human auditors to determine whether the meaning of the responses is consistent. The RMC can thus be computed as follows:
(9) $RMC=\sum_{i=1}^{N_{q}}\mathtt{Human}(\mathtt{LLM}(as_{gt}^{i},prompt_{i}),\mathtt{LLM}(as_{pred}^{i},prompt_{i}))/{N_{q}}$
where $\mathtt{Human}$ outputs 1 if the human auditor considers the LLM’s responses on ground truth and predicted action sequences share the same meaning. Specifically, for XRF V2, for the 171 test sequences, we prepared two prompts for each sequence, resulting in a total of 342 LLM responses to the ground truth and another 342 responses to the predicted results. All used prompts are presented in the Appendix. A.
We invited five auditors (3 male and 2 female), each with 1-3 years of experience using LLM models, to evaluate whether the meanings of the real and predicted responses were consistent for each of the 342 pairs. As shown in the formula, when the meanings are consistent, the numerator is incremented by 1, and the denominator represents the total number of evaluations, $N_{q}=342$. To reduce randomness, we perform action summarization using three different LLM models: ChatGPT-4o, Qwen2.5-Plus, and DeepSeek-V3, under their free plans. The final evaluation metric, mRMC, is obtained by averaging the results from different models and auditors. This provides a more robust and reliable measure for assessing the performance of action summarization.
5.2. Results
Table 4. Comparison of different methods. We evaluate several state-of-the-art approaches adapted in IMU-based temporal action localization (TAL) (Bock et al., 2024), Wi-Fi based temporal action detection (Wang et al., 2023) and TAL (Liu et al., 2025). Our method achieves the best performance, outperforming the recent WiFiTAD (Liu et al., 2025) by a large margin of 5.49 points in mAP@avg, while using 35% fewer parameters.
model mAP@0.5 0.55 0.6 0.65 0.7 0.75 0.8 0.85 0.9 0.95 mAP@avg Params GFlops
TemporalMaxer (Tang et al., 2023) 89.59 86.70 81.23 69.06 54.82 39.60 24.91 11.84 4.13 0.77 46.27 9.73M 7.93
ActionFormer (Zhang et al., 2022) 90.56 87.94 83.28 77.59 68.60 54.94 37.09 18.65 5.89 0.76 52.53 47.66M 15.70
TriDet (Shi et al., 2023) 92.90 91.34 89.51 83.37 77.91 67.50 51.46 31.19 11.61 2.28 59.91 16.12M 8.94
UWiFiAction (Wang et al., 2023) 87.59 86.32 84.12 81.94 78.33 66.77 53.11 34.23 13.17 1.84 58.74 6.24M -
WiFiTAD (Liu et al., 2025) 95.34 94.71 94.32 92.83 86.32 81.75 73.62 57.73 36.69 19.14 73.25 121.06M 44.81
ActionMamba (Chen et al., 2024) 95.05 94.84 94.10 92.81 91.54 84.63 76.56 57.54 26.89 4.35 71.83 29.35M 11.91
XRFMamba (ours) 96.67 96.23 95.68 94.53 92.02 88.66 77.92 63.96 46.53 35.13 78.74 78.38M 43.49
Table 5. mAP of 30 actions. mAP@avg for most actions remains relatively high indicating that XRFMamba performs well in XRF V2 dataset for temporal action localization.
action mAP@0.5 0.55 0.6 0.65 0.7 0.75 0.8 0.85 0.9 0.95 mAP@avg
Walk 99.65 99.33 99.33 99.33 95.20 91.55 87.82 72.56 52.50 45.95 84.32
Sit down 99.34 98.35 98.13 98.13 97.26 97.26 92.57 73.12 63.52 56.61 87.43
Stand up 99.88 99.88 99.88 99.88 94.93 91.83 81.24 68.52 48.68 41.90 82.66
Pour water into the cup 99.75 98.72 98.72 98.52 95.48 91.85 88.08 69.84 47.34 43.83 83.21
Drink water 94.99 94.99 94.99 94.99 88.97 88.97 88.64 69.76 56.24 38.10 81.07
Take medicine 98.17 98.17 98.17 98.17 96.67 92.60 85.35 75.52 62.11 46.03 85.10
Pick up things 98.73 98.73 98.73 98.73 98.73 98.73 85.18 67.18 58.82 57.18 86.07
Take the fruits from the cabinet 100.00 100.00 100.00 100.00 88.74 72.96 72.14 57.89 36.78 33.25 76.18
Cut fruits 99.75 99.75 99.75 94.87 94.29 89.80 82.64 55.46 24.33 15.75 75.64
Eat fruits 98.45 98.45 96.10 94.48 94.06 91.38 79.61 66.79 45.47 26.76 79.16
Wash hands 99.74 99.74 99.74 98.15 97.71 95.71 93.72 75.37 53.69 46.99 86.06
Throw waste 97.79 97.79 97.79 97.79 95.25 95.25 89.08 84.92 74.85 71.16 90.16
Wipe the table 97.55 94.58 92.79 89.40 89.30 89.30 83.68 83.68 72.55 59.04 85.19
Stretching 53.93 53.21 52.59 46.07 42.64 37.03 27.53 21.26 7.90 1.87 34.40
Turn on and off the desk lamp 98.42 98.42 98.42 97.46 95.54 95.54 85.38 69.30 60.77 54.82 85.41
Operate the mouse 96.43 96.43 96.43 96.43 96.43 87.42 83.49 65.83 59.69 39.77 81.84
Write 99.62 99.62 99.62 95.18 90.23 87.38 86.10 64.30 46.22 38.06 80.63
Operate the keyboard 95.97 95.97 95.97 95.97 95.97 90.98 85.58 76.38 60.35 53.76 84.69
Read a book 97.39 97.39 97.39 97.34 94.79 94.79 94.16 85.60 47.73 23.79 83.04
Open an envelope 83.14 75.93 75.07 70.76 64.49 59.00 55.48 43.17 16.88 12.14 55.61
Answer the phone 100.00 100.00 100.00 100.00 100.00 95.26 86.28 54.70 23.06 16.79 77.61
Write on the blackboard 99.43 99.43 99.43 99.43 96.40 94.29 94.29 88.82 82.64 50.75 90.49
Get up 95.42 95.42 95.42 95.42 95.42 89.25 89.25 81.90 52.41 28.52 81.84
Lie down 99.71 99.71 99.71 93.12 87.32 80.84 55.82 48.48 34.96 15.35 71.50
Use phone 99.11 99.11 99.11 99.11 99.11 91.91 85.12 35.15 19.80 5.43 73.29
Open and close windows 99.77 99.77 99.77 99.77 90.75 89.51 58.42 54.33 45.90 38.86 77.69
Open and close curtains 100.00 100.00 100.00 100.00 100.00 96.34 81.87 71.25 53.17 41.35 84.40
Water plants 99.74 99.74 99.74 99.74 97.35 97.35 88.58 69.34 51.35 26.03 82.89
Stand still 98.21 98.21 87.71 87.71 87.71 85.71 59.49 57.31 36.20 24.17 72.25
Lying still 100.00 100.00 100.00 100.00 100.00 100.00 0.00 0.00 0.00 0.00 62.22
mean 96.67 96.23 95.68 94.53 92.02 88.66 77.55 63.59 46.53 35.13 78.73
5.2.1. Temporal Action Localization Results
We evaluate several state-of-the-art approaches, including TriDet (Shi et al., 2023), ActionFormer (Zhang et al., 2022), and TemporalMaxer (Tang et al., 2023), which have been recently applied to IMU-based temporal action localization (Bock et al., 2024). We also compare with UWiFiAction (Wang et al., 2023), a WiFi-based action detection method, and WiFiTAD (Liu et al., 2025), a WiFi-based temporal action localization method. Additionally, we test ActionMamba (Chen et al., 2024), an advanced Mamba-based model for video-based action localization. All these methods are extended to support dual-input (IMU and WiFi) fusion, consistent with our approach. As shown in Table. 4, our method achieves the best performance, outperforming the recently accepted WiFiTAD (Liu et al., 2025) by a large margin of 5.49 points in mAP@avg, while using 35% fewer parameters.
We further present the mAP (mean Average Precision) for each action in Table. 5. From the table, it is clear that the mAP@avg for most actions remains relatively high, indicating that XRFMamba performs well in recognizing and localizing the majority of actions. However, some actions, such as Stretching and Opening Envelope, have relatively low mAP values compared to others. The Stretching action involves various body parts and features a wide range of motion with substantial variation in execution style. Different individuals may perform the stretching action in vastly different postures, which increases the complexity of recognizing the action accurately. In contrast, Opening an Envelope typically consists of subtle, short-duration hand movements, often involving just the fingers and palms. These localized and brief movements are more challenging to capture. Additionally, there is a sudden drop at 0.75 and 0.8 for Lying Still because it appeared only once in the test data, with its tIoUS being 0.7564.
The primary limitation of our method lies in the number of parameters that need to be learned, which is nearly 2.67 times that of ActionMamba (Chen et al., 2024). This is largely due to the use of TSSE (Temporal Signal Semantic Embedding) embedding module from WiFiTAD (Liu et al., 2025), which accounts for approximately 55.5M parameters. Therefore, a promising future research direction is to explore lighter and more efficient embedding modules that can maintain performance while significantly reducing parameters. Additionally, considering that the current state-of-the-art in video-based temporal action localization (TAL) method, such as ActionMamba (Chen et al., 2024), often relies on pre-trained models, such as InterVideo-6B (Wang et al., 2022a), as embedding modules, another potential research direction is to investigate the use of pre-trained time-series foundation models, such as TimesFM (Das et al., 2023), FoundTS (Li et al., 2024b), and Time-MoE (Shi et al., 2024), as the embedding module, thereby reducing the number of parameters that need to be learned during training.
5.2.2. Action Summarization Results
Table 6. XRFMamba achieves high response meaning consistency (RMC) in action summarization.
ChatGPT-4o Qwen2.5-Plus DeepSeek-V3 mean
RMC@a1 [278/342]0.813 [276/342]0.807 [278/342]0.813 0.811
RMC@a2 [276/342]0.807 [275/342]0.804 [276/342]0.807 0.806
RMC@a3 [279/342]0.816 [279/342]0.816 [276/342]0.807 0.813
RMC@a4 [278/342]0.813 [275/342]0.804 [277/342]0.810 0.809
RMC@a5 [262/342]0.766 [263/342]0.769 [266/342]0.778 0.771
mRMC 0.803 0.800 0.803 0.802
Figure 9. XRFMamba, when combined with LLMs, can function as an intelligent ambient sensing agent, such as personal assistant, health assistant, and home assistant, to provide services like task reminders, personalized health monitoring, climate control, and more.
We perform action summarization using three different LLM models: ChatGPT-4o, Qwen2.5-Plus, and DeepSeek-V3, all under their free plans. The consistency between predicted and real action summaries from five auditors is presented in Table 6. As shown in the table, auditors 1 to 4 consistently acknowledge that our LLMs’ predictions for TAL demonstrate high alignment with ground truth, achieving a response meaning consistency (RMC) of approximately 0.81. Auditor 5 exhibits slightly lower agreement with an RMC of 0.77. The mRMC across all auditors reaches 0.802, indicating XRFMamba achieves good TAL and action summarization results.
Notably, the auditor-specific RMC values show minimal correlation with LLM selection, indicating that XRFMamba, when integrated with free-plan LLMs, can effectively function as an intelligent ambient sensing agent. Practical implementations of such agents - including personal assistants, health monitoring systems, and smart home controllers (as illustrated in Fig. 9) - demonstrate capabilities ranging from task management and personalized health tracking to environmental regulation, thereby significantly enhancing smart home ecosystems through multifunctional intelligent services.
Fig. 10 presents prediction examples in Bedroom, Study room, and Dining room, with comparison methods including TemporalMaxer (Tang et al., 2023), ActionFormer (Zhang et al., 2022), TriDet (Shi et al., 2023), UWiFiAction (Wang et al., 2023), WiFiTAD (Liu et al., 2025), and ActionMamba (Chen et al., 2024) from Table. 4. It is visually clear that WiFiTAD, ActionMamba, and our XRFMamba perform better, with XRFMamba demonstrating superior classification accuracy and precise action boundary localization compared to state-of-the-art methods.
Figure 10. Example of the predicted results in Bedroom, Study room, and Dining room. In all the scenes, XRFMamba demonstrates excellent performance, with superior classification accuracy and precise action boundary localization compared to state-of-the-art methods.
5.2.3. Leave-one-person-out Results
In real-world applications, XRFMamba should not be limited to specific individuals but should instead be adaptable to a wide range of users with varying characteristics, such as body movement patterns, walking styles, and device usage habits. To assess this adaptability to new users, we employ a leave-one-person-out cross-validation evaluation. In this evaluation, data from one volunteer is used as the test set, while data from all other users are used for training. As shown in Table 7, mAP for new users is 72.83, indicating that our method performs well and is capable of generalizing effectively to unseen users.
In our experiments, we have not performed cross-environment testing because the actions differ across environments, as shown in Table 2. Actions in one environment may not exist in another, making cross-environment testing inconsistent and unreliable. Therefore, it is not feasible in our current setup.
Table 7. Leave-one-person-out results. mAP for new users reaches 72.83, indicating that XRFMamba performs well and is capable of generalizing effectively to unseen users.
Test user mAP@0.5 0.55 0.6 0.65 0.7 0.75 0.8 0.85 0.9 0.95 mAP@avg
user1 94.31 94.29 93.96 93.91 93.04 90.85 84.06 74.31 53.53 43.54 81.58
user2 90.08 88.89 87.84 84.72 80.64 75.15 61.94 44.69 27.68 20.13 66.18
user2 96.91 96.91 96.49 96.14 92.21 88.25 84.93 75.87 57.34 45.85 83.09
user4 96.97 96.61 95.57 94.62 90.71 85.03 76.84 61.34 41.90 25.27 76.49
user5 55.42 55.30 54.79 53.31 51.18 48.45 44.80 36.64 23.37 14.67 43.79
user6 92.63 92.16 90.86 89.57 87.91 86.37 83.59 74.51 61.00 45.90 80.45
user7 91.15 89.98 89.12 87.78 85.84 82.33 75.20 62.78 47.43 39.66 75.13
user8 90.76 90.64 90.07 89.41 84.24 78.33 71.25 58.67 39.68 31.31 72.44
user9 93.44 93.12 91.71 90.70 87.47 83.60 76.70 60.40 41.77 32.18 75.11
user10 89.21 88.21 87.14 84.68 77.40 72.43 61.59 45.25 31.87 24.79 66.26
user11 95.71 95.38 94.69 93.30 89.27 85.40 79.47 65.21 44.65 35.75 77.88
user12 92.55 92.09 90.98 89.84 85.21 78.49 71.52 56.59 35.82 25.18 71.83
user13 95.45 95.45 94.81 94.20 91.45 87.77 81.00 68.42 47.25 34.38 79.02
user14 91.23 91.22 90.65 88.31 84.83 81.50 74.48 62.79 46.23 36.23 74.75
user15 86.24 85.77 85.07 82.19 77.43 68.54 64.21 49.70 36.31 27.63 66.31
user16 93.38 92.69 91.00 88.75 85.44 81.84 73.44 60.77 47.54 35.82 75.07
Average 90.34 89.92 89.05 87.59 84.02 79.65 72.81 59.87 42.71 32.39 72.84
Table 8. We selected the weighted addition fusion method applied after the projection module (#1) as the final fusion approach for XRFMamba, as it achieves the best mAP-parameter trade-off.
# fusion fusion after mAP@0.5 0.55 0.6 0.65 0.7 0.75 0.8 0.85 0.9 0.95 mAP@avg Params Gflops
1 weight projection 96.67 96.23 95.68 94.53 92.02 88.66 77.92 63.96 46.53 35.13 78.74 78.38M 43.49
2 weight embedding 96.97 96.53 95.88 95.04 92.66 88.08 80.26 67.97 49.71 38.51 80.16 128.89M 77.58
3 weight backbone 97.12 96.83 96.57 95.88 92.91 89.80 82.69 69.48 49.42 38.19 80.89 150.08M 81.96
4 linear projection 94.87 94.04 93.34 91.83 87.63 83.03 76.53 62.03 40.87 34.26 75.84 95.16M 52.09
5 linear embedding 96.23 95.59 95.04 94.41 90.68 86.99 80.17 63.65 42.50 30.45 77.57 129.14M 77.71
6 linear backbone 95.86 95.50 94.94 90.68 89.13 83.60 75.69 59.11 34.07 08.08 72.66 150.43M 82.14
7 gate projection 96.86 96.56 95.78 91.99 89.69 86.42 79.47 65.07 45.58 36.21 78.36 79.17M 44.57
8 gate embedding 97.22 96.87 96.81 95.88 89.88 84.98 80.19 66.09 48.58 38.72 79.52 129.67M 77.85
9 gate backbone 96.92 96.69 96.41 95.58 93.27 86.03 81.01 67.56 52.83 40.93 80.72 154.03M 82.22
5.3. Ablation Study
5.3.1. How and Where to Conduct Fusion
XRFMamba takes both IMU and WiFi modalities from XRF V2 as input. We first test where and how modality fusion should be applied in the XRFMamba pipeline. We evaluate fusion after the projection, after embedding, and after the Mamba backbone. In terms of fusion strategies, we evaluate three methods: (1) simple weighted addition: applying fixed weight coefficients to the IMU (0.8) and Wi-Fi (0.2) features and directly adding the weighted features; (2) linear fusion: combining the IMU and Wi-Fi features through a learned linear transformation, followed by a gating mechanism using a sigmoid function to weight the contribution of each modality; and (3) gate fusion: applying separated convolutional layers to the IMU and Wi-Fi features to compute individual gating weights using a sigmoid function, which are then used to weight each modality and combine them through summation. For a detailed explanation of each fusion method, including the mathematical formulations and pseudocode, please refer to the Appendix. B.
The results of the nine different combinations of fusion positions and strategies are shown in Table 8. As illustrated in the table, fusion applied later in the pipeline involves more separate data streams and additional parameters. When comparing different fusion strategies, the gate and linear methods require more parameters than the weighted method, yet their performance is not as effective. In terms of fusion positions, combination 1, which applies weighted fusion after the projection module, achieves an average mAP of 78.74. In contrast, combination 3, which performs fusion after the Mamba backbone, achieves a higher mAP of 80.89 but requires nearly double the number of parameters. Therefore, we chose the weighted addition fusion method applied after the projection module as the final fusion approach for XRFMamba, as it offers the best mAP-parameter trade-off.
device mAP@0.5 0.55 0.6 0.65 0.7 0.75 0.8 0.85 0.9 0.95 mAP@avg
1 95.88 95.32 93.52 90.90 85.02 76.06 60.50 40.57 17.40 2.93 65.81
2 94.07 93.03 91.17 87.33 79.99 67.60 51.39 27.82 10.97 1.75 60.51
3 80.99 79.94 77.65 74.81 71.11 64.19 53.78 37.90 19.02 2.86 56.22
4 91.91 90.38 85.54 79.48 68.81 54.26 35.51 18.48 7.98 1.16 53.35
5 85.76 83.82 80.99 76.34 70.48 56.36 43.89 29.39 13.03 2.76 54.28
6 89.06 88.14 86.48 83.64 75.66 66.33 50.27 30.88 9.77 1.42 58.16
7 65.40 60.51 54.67 47.75 34.70 24.68 13.17 5.52 1.67 0.32 30.84
8 95.58 95.07 93.77 88.44 85.46 77.62 67.22 49.08 26.28 4.91 68.34
9 95.19 94.60 92.87 90.86 86.38 78.39 65.48 46.95 18.51 4.58 67.38
10 91.03 89.70 87.84 82.18 78.23 71.02 57.76 37.97 18.55 4.59 61.89
11 92.40 91.80 90.77 87.64 84.49 76.19 65.69 49.87 28.32 7.66 67.48
12 94.86 94.02 93.50 91.69 89.33 82.84 71.61 52.37 22.11 3.06 69.54
13 95.29 94.60 93.57 92.19 89.82 83.28 68.06 44.47 18.34 2.76 68.24
14 94.36 94.15 92.62 90.66 87.59 80.33 68.98 51.75 31.19 10.64 70.23
15 93.05 92.76 91.91 90.15 86.76 81.01 69.77 54.25 37.55 17.17 71.44
16 96.22 95.16 94.43 92.24 89.17 79.57 68.14 47.16 22.77 4.58 68.94
17 96.36 95.87 95.26 93.23 90.65 87.08 75.10 59.85 38.74 11.36 74.35
18 96.82 96.51 96.23 95.31 92.84 84.06 73.08 54.41 26.99 3.95 72.02
19 96.51 95.82 95.44 94.23 92.00 86.98 80.26 62.57 41.40 29.00 77.42
20 97.48 97.33 96.77 92.02 89.02 84.29 76.69 60.04 41.98 18.64 75.43
21 96.67 96.23 95.68 94.53 92.02 88.66 77.92 63.96 46.53 35.13 78.74
Figure 11. The XRF V2 dataset allows combinations of different devices, and we tested the most common device combinations, including individual devices and their configurations.
5.3.2. Device Combination Study
XRF V2 dataset includes Wi-Fi signals and IMU data from seven types of devices: phones, watches, earbuds, and glasses, as shown in Fig.11. Users can combine devices according to their own needs. In Fig. 11, we report some of the most common device combinations, such as case 9 (a phone and a watch), case 14 (a phone, a watch, and an earbud), case 17 (a phone, a watch, glasses, and ambient Wi-Fi). It shows that the average mAP of using only a right-hand-worn smartwatch (case 1) is 65.81, and case 17 achieves an average mAP of 74.35, reaching a usable level.
6. Dataset and Code Availability, and Potential Usefulness
The data and code are available at https://github.com/aiotgroup/XRFV2. The repository includes a comprehensive $\mathtt{readme.md}$ file, which describes how to use the code to reproduce our results. The code features a well-structured modular design, making it easy to call and test various backbones and modality combinations. In this code repository, under the /model directory, you will find different method implementations, including XRFMamba and all comparison methods. You can easily use the model you need through make_model() found in /model/models.py. The code for XRFMamba is located in model/TAD_multi_fusion.py. Data loading methods are defined in the /dataset directory. To get started, you need to modify the paths in basic_config.json to match your system setup. Then, run script/train_run.py to train the model. Finally, you can evaluate the model by running script/test_run.py. For more details, please check out our repository.
We believe our dataset can significantly contribute to the community in at least the following areas.
  1. (1)
    The XRF V2 dataset can serve as a benchmark for action recognition, temporal action localization, and action summarization based on smartphones, watches, earbuds, glasses, and Wi-Fi. It supports research aimed at developing methods that either achieve higher precision or are more lightweight.
  2. (2)
    XRF V2 also supports weakly supervised action localization, where only the action categories are provided without specifying the precise start and end times of the actions. This setup allows for research on action localization methods that can infer temporal boundaries from less detailed supervision, enabling more scalable and practical approaches for real-world applications.
  3. (3)
    XRF V2 is a dataset collected using real devices, and it provides high-quality data labels. It can be used to promote or evaluate research on IMU and WiFi data synthesis methods.
  4. (4)
    XRF V2 supports the research of next action prediction, which facilitates the development of systems that can anticipate and respond to user actions. For example, by detecting the user’s intent to turn on the lights, the system can automatically trigger the action without the need for voice or manual control. This capability enhances the intelligence of smart devices, making them more context-aware and responsive.
  5. (5)
    XRF V2 supports multimodal learning between IMU, WiFi, and video, facilitating the development of multimodal foundational models. While the original videos cannot be shared due to privacy concerns with volunteers, we are to process them into features using InterVideo-6B (Wang et al., 2022a) and make these features publicly available.
  6. (6)
    We are to process the videos into 2D pose, 3D pose, and human mesh, which will support research in pose estimation or mesh reconstruction based on WiFi, IMU, and video data. Compared to previous datasets such as Person-in-WiFi 3D (Yan et al., 2024) and MM-Fi (Yang et al., 2024), where the person stands at a fixed position and repeats actions, the actions in XRF V2 are more representative of real-world scenarios, capturing more dynamic and varied human behavior.
  7. (7)
    The academic and industrial sectors are welcome to leverage our dataset to develop specific applications, such as ambient sensing systems and context-aware systems, which help in enhancing smart home automation, health, and wellness monitoring.
7. Conclusion
We present the XRF V2 dataset, which captures continuous action sequences from 16 volunteers across three typical household environments: the dining room, study room, and bedroom. Comprising a total of 825 action sequences, the dataset integrates data from various sensors, including IMUs from smartwatches, smartphones, earbuds, and glasses, Wi-Fi transceivers, and synchronized video from an Azure Kinect. The diverse and personalized action sequences, along with synchronized multimodal data, provide a valuable resource for advancing research in human action understanding and multimodal machine learning. We also introduce a novel human action understanding task: action summarization, for which we design a new evaluation metric, response meaning consistency (RMC). Action summarization takes action sequences and task-oriented prompts as input to large language models (LLMs), transforming the LLMs into intelligent agents. We demonstrate that with IMU or Wi-Fi input, LLMs can function as personal assistants, health assistants, and home assistants. Furthermore, we propose the XRFMamba neural network, which effectively performs temporal action summarization and action summarization, and conduct a detailed evaluation showing its superior performance compared to existing state-of-the-art methods.
Appendix A Evaluated Task-orientated Prompts
A.1. An example of responses from three LLMs to the same prompt.
—— Input:#### You are an intelligent action summarization agent. #### This is an action sequence that occurred in the bedroom: the user did the action of walking, with a start time of 0.0 and an end time of 5.0; the user did the action of sitting down, with a start time of 5.0 and an end time of 13.0; the user did the action of reading a book, with a start time of 13.0 and an end time of 22.0; the user did the action of pouring water into the cup, with a start time of 22.0 and an end time of 27.0; the user did the action of taking medicine, with a start time of 27.0 and an end time of 40.0; the user did the action of lying, with a start time of 40.0 and an end time of 46.0; the user did the action of using the phone, with a start time of 46.0 and an end time of 60.0; the user did the action of getting up, with a start time of 60.0 and an end time of 68.0. #### Please answer whether the user read a book before lying down during this period?”
——Output of ChatGPT-4o: “Yes, the user read a book before lying down during this period. The action of reading a book occurred from 13 to 22. The action of lying down occurred from 40 to 46. Therefore, reading the book happened before lying down.”
——Output of Qwen2.5-Plus: “Yes, the user read a book before lying down during this period. According to the sequence of actions: The user started reading a book at 13 and ended at 22. The user then performed other actions (pouring water, taking medicine) before starting to lie down at 40. Since the action of reading a book occurred before the action of lying down, the answer is yes, the user did read a book before lying down.”
——Output of DeepSeek-V3: “Yes, the user read a book before lying down during this period. The action of reading a book occurred from 13 to 22, and the action of lying down occurred from 40 to 46. Therefore, the user read a book before lying down.”
A.2. All task-oriented questions used in action summarization evaluation.
(1) Please answer the number of times the user has consumed water during this period?
(2) Please answer whether the user has been drinking water during this period?
(3) Please answer whether the user has taken medication during this period?
(4) Please answer the number of times the user has taken medication during this period?
(5) Please answer whether the user has been reading during this period?
(6) Please answer whether the user read a book before lying down during this period?
(7) Please answer the number of times the user watered the plants during this period?
(8) Please answer whether the user watered the plants during this period?
(9) Please answer whether the user has opened windows for ventilation during this period?
(10) Please answer whether the user was playing with their phone while lying in bed during this period?
(11) Please answer whether the user has been walking during this period?
(12) Please answer the number of times the user has stretched during this period?
(13) Please answer whether the user has eaten fruits during this period?
(14) Please answer the number of times the user has eaten fruits during this period?
(15) Please answer whether the user has wiped the table during this period?
(16) Please answer whether the user has thrown away garbage during this period?
(17) Please answer whether the user washed their hands before eating fruits during this period?
(18) Please answer whether the user washed their hands after eating fruits during this period?
(19) Please answer whether the user has washed their hands after littering during this period?
(20) Please answer whether the user has washed their hands after wiping the table during this period?
(21) Please answer how many times the user washed their hands during this period?
(22) Please answer whether the user operated the mouse during this period?
(23) Please answer whether the user operated the keyboard during this period?
(24) Please answer whether the user has opened the envelope during this period?
(25) Please answer whether the user has turned on the desk lamp during this period?
(26) Please answer whether the user answered the phone during this period?
(27) Please answer the number of times the user answered the phone during this period?
Appendix B Detailed Description of Three Fusion Methods
In this section, we describe three different modality fusion strategies: Gated Fusion, Linear Fusion, and Weighted Fusion. Each method combines IMU and Wi-Fi features using different techniques. Below, we provide both the mathematical formulation and pseudocode for each method.
B.1. Weighted Fusion
In Weighted Fusion, each modality is given a fixed weight, and the features are combined by weighted addition. The IMU features are weighted by $w_{\text{IMU}}=0.8$ and the Wi-Fi features by $w_{\text{WiFi}}=0.2$ before being summed.
Mathematical Formulation:
$\mathbf{X}_{\text{IMU}}^{\text{weighted}}=w_{\text{IMU}}\times\mathbf{X}_{\text{IMU}}$
$\mathbf{X}_{\text{WiFi}}^{\text{weighted}}=w_{\text{WiFi}}\times\mathbf{X}_{\text{WiFi}}$
$\mathbf{X}_{\text{fused}}=\mathbf{X}_{\text{IMU}}^{\text{weighted}}+\mathbf{X}_{\text{WiFi}}^{\text{weighted}}$
Algorithm 2 Weighted Fusion Implementation
def weighted_fusion(imu_features, wifi_features, w_imu=0.8, w_wifi=0.2):
# Apply fixed weights to IMU and WiFi features
gated_imu = w_imu * imu_features
gated_wifi = w_wifi * wifi_features
# Combine the weighted features by addition
fused_features = gated_imu + gated_wifi
return fused_features
B.2. Linear Fusion
Linear Fusion applies a learned linear transformation to combine IMU and Wi-Fi features, followed by a gating mechanism. The resulting features are combined using the learned gate and passed through a fully connected layer.
Mathematical Formulation:
$\mathbf{G}=\sigma\left(\mathbf{W}_{g}\cdot\left(\mathbf{X}_{\text{IMU}}+\mathbf{X}_{\text{WiFi}}\right)\right)$
$\mathbf{X}_{\text{fused}}=\text{fc}\left(\mathbf{G}\times\left(\mathbf{X}_{\text{IMU}}\oplus\mathbf{X}_{\text{WiFi}}\right)\right)$
where $\oplus$ denotes concatenation.
Algorithm 3 Linear Fusion Implementation
def linear_fusion(imu_features, wifi_features):
# Combine IMU and WiFi features element-wise
combined = imu_features + wifi_features
# Apply a linear transformation to the combined features
gate = sigmoid(linear_transformation(combined))
# Concatenate the IMU and WiFi features
concatenated_features = torch.cat([imu_features, wifi_features], dim=-1)
# Apply gating mechanism (element-wise multiplication)
gated_features = gate * concatenated_features
# Pass the gated features through a fully connected layer
fused_features = fully_connected(gated_features)
return fused_features
B.3. Gated Fusion
Gated Fusion uses a learned gating mechanism to weight the contribution of each modality. A 1D convolution is applied to the IMU and Wi-Fi features, followed by a sigmoid activation function to compute gating weights. These gated features are then combined element-wise.
$\mathbf{G}_{\text{IMU}}=\sigma\left(\mathbf{W}_{g}\cdot\mathbf{X}_{\text{IMU}}\right),\quad\mathbf{G}_{\text{WiFi}}=\sigma\left(\mathbf{W}_{g}\cdot\mathbf{X}_{\text{WiFi}}\right)$
$\mathbf{X}_{\text{IMU}}^{\text{gated}}=\mathbf{G}_{\text{IMU}}\times\mathbf{X}_{\text{IMU}},\quad\mathbf{X}_{\text{WiFi}}^{\text{gated}}=\mathbf{G}_{\text{WiFi}}\times\mathbf{X}_{\text{WiFi}}$
$\mathbf{X}_{\text{fused}}=\mathbf{X}_{\text{IMU}}^{\text{gated}}+\mathbf{X}_{\text{WiFi}}^{\text{gated}}$
Algorithm 4 Gated Fusion Implementation
def gated_fusion(imu_features, wifi_features):
# Apply 1D convolution to IMU and WiFi features
gate_imu = sigmoid(convolution(imu_features))
gate_wifi = sigmoid(convolution(wifi_features))
# Apply gating mechanism (element-wise multiplication)
gated_imu = gate_imu * imu_features
gated_wifi = gate_wifi * wifi_features
# Combine the gated features by addition
fused_features = gated_imu + gated_wifi
return fused_features